OpenStudy (anonymous):

Solve 5 sin x - 3 cos x = 0, for 0 ≤ x ≤ 360 degrees.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

tanx=3/5 x=arectan.6

OpenStudy (anonymous):

x=30.94+npi

OpenStudy (anonymous):

Thank you!